### v3

# Combining Threat Intelligence with IoT Scanning to Predict Cyber Attacks

Abstract
While the Web has become a global platform for communication; malicious actors, including hackers and hacktivist groups, often disseminate ideological content and coordinate activities through the ”Dark Web”—an obscure counterpart of the conventional web. Presently, challenges such as information overload and the fragmented nature of cyber threat data impede comprehensive profiling of these actors, thereby limiting the efficacy of predictive analyses of their online activities. Concurrently, the proliferation of internet-connected devices has surpassed the global human population, with this disparity projected to widen as the Internet of Things (IoT) expands. Technical communities are actively advancing IoT-related research to address its growing societal integration. This paper proposes a novel predictive threat intelligence framework designed to systematically collect, analyze, and visualize Dark Web data to identify malicious websites and correlate this information with potential IoT vulnerabilities. The methodology integrates automated data harvesting, analytical techniques, and visual mapping tools, while also examining vulnerabilities in IoT devices to assess exploitability. By bridging gaps in cybersecurity research, this study aims to enhance predictive threat modeling and inform policy development, thereby contributing to intelligence research initiatives focused on mitigating cyber risks in an increasingly interconnected digital ecosystem.
keywords:
Internet of Things, Dark Web, IoT Vulnerability, Machine Learning, Cybersecurity, Threat Intelligence
1 Introduction
The Internet has evolved into a worldwide platform that allows individuals to easily distribute, exchange, and communicate ideas. While it offers numerous benefits, the misuse of the Internet has grown increasingly problematic. Malicious entities, including hackers, extremist organizations, hate groups, and racial supremacists, are leveraging the web more than ever to carry out illegal activities, facilitate internal communications, orchestrate attacks on institutions and governments, and spread harmful ideologies. It has become commonplace to hear about cyber breaches in the news, where cybercriminals post messages targeting large corporations and organizations, demanding weapons, financial backing, and volunteers [1]. Consequently, there is a pressing need to gather intelligence that enhances our understanding of the tactics employed by hackers and criminal organizations.
In this paper, I focus on the Internet of Things (IoT) as a vulnerability due to its expanding presence and susceptibility to exploitation [2]. By employing the structured methodology outlined in this paper, I evaluate the degree to which cybercriminal groups target IoT systems. Through my analysis of 23 criminal websites on the Dark Web, I discovered that even seemingly harmless devices, such as small sensors, can cause substantial damage when compromised. IoT opens up an entirely new realm for cybercriminals to exploit. These hacker groups generally do not discriminate in their targets—they exploit any weakness or vulnerability they can find.
The remainder of the paper is organized as follows: the second section provides an overview of the concepts introduced in this paper; the third section outlines a methodology for collecting, analyzing Dark Web data, and identifying vulnerabilities[3]. The fourth section details how this methodology was applied in a case study of IoT using Shodan, the IoT search engine. The final section concludes the study and explores potential solutions.
2 Literature Review
2.1 Dark Web Utilization
Hackers, criminals, and terrorists leverage the web to facilitate their illicit activities. Since the emergence of the Dark Web, numerous new marketplaces have emerged, offering a wide and often alarming array of products, including prescription pills, meth, heroin, speed, crack, guns, stolen identities, gold, and erotica. The Dark Web is also a platform for spreading extremism, facilitating communication among terrorists and criminals, and sharing knowledge about illegal hacks and attacks, thereby enabling harmful activities that threaten society. Most of these markets operate using Bitcoin and other cryptocurrencies that allows for anonymous online transactions[4]. According to one study, criminals and terrorists use Bitcoin to sell drugs and subsequently purchase guns and ammunition. During my exploration of the dark market known as Hansa Market (http://hansamkt2rr6nfg3.onion/search/?q=hacker&c=59), I identified a website called Arms and Ammunition (http://armsforsd.com/index.html), which poses significant threats to nations both internally and externally. This underscores the importance of developing robust threat intelligence.
2.2 Discovering IoT Devices Through Shodan
I chose to focus on IoT devices due to their profound impact on various aspects of daily life. These devices, similar to our mobile phones, will become integral to our routines and will predict our behavior. Despite their numerous applications, IoT devices are highly susceptible to hacks and malfunctions. Many of these devices operate on low-power CPUs and custom operating systems that lack the capability to support encryption and decryption mechanisms, making them inherently insecure. For instance, a smart thermostat does more than just regulate temperature; it stores data on a central server that helps predict user temperature patterns. While a compromised thermostat in a home may not be immediately dangerous, in industrial control systems, even minor data manipulation through sensors can have catastrophic consequences. If hackers or criminals exploit the vulnerabilities of these critical IoT devices, they could potentially threaten individuals based on their past behavior or disrupt essential services like power stations[5]. To identify vulnerable IoT devices, I utilized Shodan. Shodan is an IoT search engine that scans the internet to locate devices with open ports and establish successful connections. It can identify a wide range of internet-connected devices, including computer servers, webcams, industrial control systems, power plants, traffic signals, and virtually any other device. Shodan is a relatively new search engine specifically designed for IoT devices, and its reliability has been validated by several research studies. One such study, conducted over several months, connected a few devices to the internet for the first time to see if Shodan would list them. After 14 days, the study found that the devices’ information, including IP addresses, ports, locations, and device details, appeared on Shodan [6]. As per findings, Shodan collects this information through service banner interrogation, which involves extracting data from a system in response to a connection request. Banner grabbing is the process of obtaining active information about a service or system through port interrogation. With various filters available - such as IP address, port, country, city, latitude, longitude, and device information (OS) - one can identify and exploit vulnerable devices[7]. However, this capability is invaluable for developing threat intelligence.
2.3 Developing Cyber Threat Intelligence
My research focused on determining whether the systematic collection and processing of content from the Dark Web could contribute to the development of actionable threat intelligence for protecting vulnerable IoT devices. I specifically examined how insights gathered from underground hacker forums, marketplaces, and other illicit platforms could enhance the understanding of emerging cyber threats targeting IoT ecosystems[8, 9]. The remainder of this paper details the methodologies I employed for data collection, the techniques I used to filter and refine relevant information, and the analytical approaches I applied to transform raw data into meaningful threat intelligence. By outlining these processes, this study demonstrates how Dark Web intelligence could have been leveraged to strengthen IoT security measures and mitigate cyber risks before they materialized into active threats.
3 Methodology
3.1 Data Collection Framework
To address the diverse threats posed by the information sources that hackers, criminals, and terrorists use to conduct destructive activities, I propose a comprehensive threat intelligence framework. This framework integrates Dark Web data aggregation, machine learning-driven anomaly detection for IoT vulnerability patterns, and human-AI collaboration for real-time cyber-attack prediction.
Figure 1: A methodology for collecting and analyzing Dark Web information.
Unlike previous methodologies [10], this approach emphasizes the correlation of hacker discussions (e.g., IoT exploitation trends) with Shodan-derived device vulnerabilities to prioritize high-risk targets. Figure 1 illustrates the methodology, which is designed to effectively assist human investigators in obtaining Dark Web intelligence through various information sources, collection methods, filtering, and analysis [11].
  • Information Sources encompass a broad spectrum of providers of hacker-related information on the web. These sources range from readily accessible platforms, such as search engines, to websites developed and maintained by hacker groups, hacktivists, terrorist organizations[12], or their supporters. Access to the latter often requires the assistance of domain experts.
  • Collection methods enable the systematic searching and browsing of information from identified sources[13]. IoT Threat Horizon Scraping begins with a set of relevant seed URLs or keywords, which are derived from the initial information gathered through browsing. IoT threat scraping can be done using various web crawling tools.
  • Back-link search, supported by search engines such as Google (www.google.com) and AltaVista (www.altavista.com, which was acquired by Yahoo! in 2003), allows investigators to find web pages that link to a target web domain or page. This feature is particularly useful for tracing the activities of hacker supporters and sympathizers, whose web pages often reference hacker sites[14]. For example, these pages may glorify hacking actions or demonstrate a pattern of coordinated attacks.
  • Personal/People Data Search, exemplified by major web portals such as Yahoo! (members.yahoo.com) and MSN (groups.msn.com), uncovers the profiles of groups or individuals with shared interests. Hackers and their supporters may include ”hot links” in their profiles[15], which enable investigators to uncover hidden connections.
  • Meta-searching leverages related keywords as input to query multiple search engines, allowing investigators or automated programs to collate top-ranked results and eliminate duplicates to obtain highly relevant information. For example, searching for the hacker group ”Anonymous” (widely known and famous for attacking the FBI, CIA, PayPal, Sony, MasterCard, Visa, Chinese, and Israeli governments) across multiple search engines yields mixed results about their hacking news articles and websites. Augmenting ”Anonymous” with terms like ”digital Robin Hood,” ”freedom fighters,” or ”Operation Anons” (as glorified by their supporters) can yield more targeted and relevant results[16].
  • Filtering involves sorting through collected information and eliminating irrelevant results. This task requires domain knowledge and linguistic expertise[17]. Domain knowledge encompasses understanding hackers, terrorists, and their supporter groups, including their presence on and usage of the web, as well as their histories, activities, and missions. Linguistic knowledge pertains to recognizing terms, slogans, and other textual and symbolic clues in the native language of cyber-attackers.
  • Analysis provides insights into data and aids investigators in identifying trends and verifying conjectures. Several functions support these analytical tasks. IoT Exploit Motive Profiling identifies common properties among entities and categorizes them to help investigators predict trends[18]. Clustering organizes entities into naturally occurring groups, facilitating the identification of similar hacker groups. IoT Risk Topography Mapping presents data in a visually comprehensible format, enabling investigators to visualize relationships within the networked structure of these hacker groups or their activities. Threat Scoring ranks vulnerabilities based on hacker interest and entity exposure. Vulnerability Discovery presents methods for selecting an entity and uncovering various threats that could be exploited for that chosen entity.
3.2 Implementation Approach
Although the Internet has been publicly accessible since the 1990s, the Dark Web has only emerged as a significant domain in recent years. Its anonymous and decentralized nature has facilitated a range of illicit activities, yet the development of effective methodologies for identifying vulnerabilities and tracking cybercriminal activities within this hidden space has remained a persistent challenge [19]. The absence of robust and scalable techniques has hindered law enforcement agencies, cybersecurity researchers, and policymakers in their efforts to combat cybercrimes effectively.
As discussed earlier, the proposed methodology integrates a combination of advanced data mining, web mining, and machine learning techniques while leveraging human domain expertise to refine their application. This hybrid approach enhances the ability to systematically identify and monitor criminal groups, analyze their discussions, and extract intelligence relevant to emerging cyber threats[20]. By incorporating both automated data processing and human analytical oversight, this methodology balances computational efficiency with contextual accuracy, enabling more precise intelligence gathering[4]. Ultimately, this semi-automated framework presents a scalable solution for uncovering hidden threat actors and understanding the evolving landscape of cybercrime on the Dark Web.
4 Case Study: IoT Scanning to Predict Cyber Attacks
To illustrate the effectiveness of the proposed methodology, we applied it to collect and analyze web-based activities of hackers, terrorists, and criminals targeting IoT, a technology with inherent security vulnerabilities. Initially, we identified dark-market sites, forums, and blogs, followed by an analysis of IoT exploitation through these platforms. Research suggests that nearly 57% of IoT devices are vulnerable [17], despite representing just 0.1% of all internet-connected devices. By 2030, projections indicate that over 26 billion IoT devices will be online, expanding the attack surface for cybercriminals. Given these trends, our case study focused on IoT security threats.
4.1 Collection
To gather data, I first identified several suspicious URLs using web searches, referencing dark web markets[21, 22]. I conducted web spidering using the ’Open Web Spidering’ tool. For instance, starting with keywords like ’freedom fighters’ and ’digital Robin Hood,’ I found four intriguing URLs: HackHound (a forum for hackers), Du Bist Anonymous (a website supporting the Anonymous group), CiHad.net (a suspicious website associated with a terror group), and Intel Exchange (a political ideology website). Using these categories as references, we uncovered additional websites during data collection. I then performed the same search on dark web search engines such as Torch, Grams, and Ahmia, although the results were scattered. I stored them nonetheless, as they primarily pointed to dark web markets. Manually browsing these markets revealed offers of both hacking services and services required for hacking. I selected keywords such as ’hacking services,’ ’free world,’ ’personal army,’ and ’Internet of Things,’ yielding various results. Overall, I collected information from hundreds of sources, and it would be an interesting future direction to study how data collection techniques affect the quality of analysis.
4.2 Filtering
I conducted two rounds of filtering. First, I manually removed unrelated sites, such as news, games, and porn sites, which had no reference to hacking. I retained websites of forums, blogs, ideological discussions, dark markets for hacking and ammunition services. After this round, thirty-two sites remained. Second, given my limited domain expertise and knowledge of languages, I initially manually removed some websites that felt less relevant and similar to English forum websites in my results. Subsequently, I used natural language processing (NLP) and machine learning (ML) for keyword tagging and threat scoring.
4.3 Analysis
4.3.1 Classification and Usage of Websites
I performed clustering and classification of the remaining 23 websites. These websites provided a diverse range of information necessary for my analysis. Table 1 lists these websites with hyperlinked URLs and descriptions.
Website Description
Hansa Market Dark market - Hacking Services
Arms and Ammunition Guns & Ammunition
Du Bist Anonymous Hacking services (Anonymous Hackers Site)
Rent A Hacker Hacking services
Alpha7-Bravo- Blog Political Hate Ideology
Thoman Paine Common Sense Political Hate Ideology
Random United Kingdom Hate group forum
Intel Exchange Hacking Forum
Bugged Planet Hate group forum
Jaggers Blog Hacking Blog and Forum
TorChan Hacking Discussions Forum
HackerWeb Cyber Security Hacking Discussion Forum
HackHound Cyber Security Hacking Discussion Forum
HackersTribe Hacking Discussions Forum
School-of-HackNet Hacking Discussions Forum
iC0de Cyber Security Hacking Discussion Forum
Valhalla Dark market - Hacking Services
Rhyliv Cyber Security Hacking Discussion Forum
Alokbab Suspicious Terror Group Site
Fadaian Suspicious Terror Group Site
CiHad.net Bogus Site – Suspicious Terror Group
Oasis Dark market - Hacking Services
Table 1: Websites collected with the implementation of the methodology.
I followed following steps to analyze “IoT Hacking” or “hacking IoT devices” discussions on some of these websites: Filtered Cyber Security Hacking Discussion Forums from above list ignoring rest of the websites from my list. Manually browsed through these forums and searched for above keywords in these forums. Gathered the statistics of as recent (to 2024) as possible and generated a graph, shown in Figure 2, and a table shown in Table 2. The graph in Figure 2 displays the percentage of IoT hacking-related discussions occurring in the seven forums I chose for analysis. The x-axis shows the names of these forums, while the y-axis shows the total percentage of IoT hacking discussions. The ”Hacker Web” forum, which is a collection of 18 major forums, has 3.3
Figure 2: Statistics of IoT device hacking discussions within December 2023 - July 2024.
Website Searched Found # of Posts
HackHound Internet of Things Botnet, Malware, Rats for Android devices 4
Hackers Tribe Hack Internet of Things Sensors, File Sharing in devices 5
School-of-HackNet Hack devices Phones Listening to Network 1
HackerWeb Hack Internet of Things Knife for IoT: XMPP 1
Table 2: This table shows the keywords searched and data found on the mentioned websites.
By manually browsing through these forums, I analyzed the most frequently discussed topics related to IoT hacking. I selected four websites listed in Table 2, searched for the keywords in the ’Searched’ column, and recorded the results and total post counts in the ’Found’ and ’# of Posts’ columns, respectively. My analysis revealed that botnets, malware, and sensors are the most discussed methods for exploiting IoT devices. To determine how easily these vulnerabilities can be exploited, I used the IoT search engine, Shodan.
4.3.2 Shodan: Discovering Vulnerable IoT Devices
Using Shodan, I followed these steps for my analysis:
  1. 1.
    Configured Shodan’s REST API with Python.
  2. 2.
    Searched for the keyword sensor.
  3. 3.
    Exported the results in .csv format.
  4. 4.
    Used Excel for further analysis.
Listing 1: Python code used for searching in Shodan and saving results to CSV
1# Importing the necessary libraries
2import shodan
3import csv
4
5# Define your Shodan API key
6SHODAN_API_KEY = "your_api_key"
7
8# Initialize the Shodan API client
9api = shodan.Shodan(SHODAN_API_KEY)
10
11# Open a CSV file to store the results
12with open(’shodan_results.csv’, ’w’, newline=’’) as csvfile:
13 fieldnames = [’IP’, ’Data’]
14 writer = csv.DictWriter(csvfile, fieldnames=fieldnames)
15
16 # Write the header to the CSV file
17 writer.writeheader()
18
19 try:
20 # Perform a search query using the Shodan API
21 results = api.search(’sensor’)
22
23 # Print the total number of results found
24 print(’Resultsfound:{}’.format(results[’total’]))
25
26 # Loop through each result and save it to the CSV
27 for result in results[’matches’]:
28 print(’IP:{}’.format(result[’ip_str’]))
29 print(result[’data’])
30
31 # Write each result to the CSV file
32 writer.writerow({’IP’: result[’ip_str’], ’Data’: result[’data’]})
33
34 # Handle API-related errors
35 except shodan.APIError as e:
36 print(’Error:{}’.format(e))
Figure 3: Example of vulnerable temperature-sensing device found through Shodan
Figure 4: Example of vulnerable sensor devices found through Shodan
Figure 5: Example of vulnerable pressure-sensor device found through Shodan
Figure 6: Example of full access admin control panel of pressure-sensing device found through Shodan
Upon executing the Python script detailed in Listing 1, I obtained 582 results for sensors. In the try block of the code, I captured the number of results and filtered information such as company name, IP address, port, operating system, and data details, including content type, authentication protocol, and location specifics like country and city. The results from my Shodan search provided comprehensive information about the devices, including company name, location, IP address, and port. This level of detail facilitates various forms of analysis. First, I filtered the results to include only those with port 8080 and directly accessed some of the IP addresses. As shown in the Figure 3, I encountered a page requesting login credentials. This is concerning because it exposes the device to a range of security vulnerabilities, such as cross-site scripting, denial-of-service attacks, SQL injection, unexpected application termination, and arbitrary code execution. Additionally, the connection was not secured over SSL. Second, I manually examined the IP addresses providing direct access. As illustrated in Figure 4 and Figure 5, I gained access to a pressure-sensing device named ’Pbar-box4’ and a temperature-sensing device named ’Byt.’ I had full control over these devices, allowing me to view past temperature readings and manipulate administrative settings, including changing email addresses where data was being sent. Figure 6 demonstrates an example of the control panel for the ’Byt’ device.
5 Conclusions
Analyzing and collecting Dark Web information is exceedingly challenging due to cyber-attackers’ ability to conceal their identities and erase traces of their activities online. The sheer volume of web information complicates the task of obtaining a comprehensive understanding of their operations, making it even more difficult to predict their next targets. In this article, I propose a methodology to address these challenges. By employing web mining, content analysis, and visualization techniques, the methodology leverages various information sources and analyzes 23 hacker websites for IoT vulnerabilities. Information clustering is used to differentiate between novice and experienced hackers on the web. My detailed evaluation indicates that the methodology produced promising results, which could significantly aid in discovering new vulnerabilities and their sources.
For IoT, I propose the following solutions:
  1. 1.
    Strict security standards must be established for the manufacturing of IoT hardware and software.
  2. 2.
    IoT software should be developed by dedicated companies that rigorously adhere to IoT security protocols.
  3. 3.
    Best practices must be followed by users and those configuring IoT devices.
This research was conducted from multiple angles, and I believe it will contribute to the existing literature on cyber-security, potentially guiding policy-making and threat intelligence research.
Statements and Declarations
Competing Interests
The author declares no competing interests.
Data Availability
From the datasets analyzed during this study include, a sample of 50 rows from Shodan scan results and 50 rows from Dark Web forum metadata is provided to demonstrate the methodology. These datasets do not contain real IP addresses, device identifiers, or sensitive information. The full datasets are available from the corresponding author upon reasonable request. Figures 1 (methodology flowchart), 2 (forum discussion statistics), and 4 (anonymized IoT device screenshots) are included in this manuscript, with all sensitive details redacted to prevent misuse.
5.1 Funding
Not applicable.